OpenStudy (anonymous):

What is the solution of the equation? 16 = –d + 6

OpenStudy (karatechopper):

subtract the 6 on both sides 16=-d+6 16-6=-d-6 -------- 10=-d divide negative on both sides 10=-d 10/-1= -d/-1 ------------ -10=d

OpenStudy (anonymous):

TY:)